OpenStudy (anonymous):

6. A 58 meter garden hose is stretched diagonally across a yard. If the width of the yard is 24 meters, what is the length of the yard, to the nearest tenth?

OpenStudy (anonymous):

ik this is english if that is what ur gonna say

OpenStudy (anonymous):

go yo math and it will be faster ♥♥♥ can i plzzz have a medal

OpenStudy (anonymous):

o.o use the Pythagorean theorem

OpenStudy (anonymous):

no medal

OpenStudy (anonymous):

This is the wrong section, but if you go to mathematics they can help you. Im more of an English person.

OpenStudy (anonymous):

ok, I will post it there!

OpenStudy (anonymous):

no one asked

OpenStudy (anonymous):

wow u r RUDE

OpenStudy (anonymous):

\[a ^{2} + b ^{2} = c ^{2}\]

OpenStudy (anonymous):

Im rude

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i want a medal

OpenStudy (anonymous):

~sighs~ o-o dude calling someone rude is infact rude.

OpenStudy (anonymous):

why should I give you a medal ur not helping me @BVB4rmy is

OpenStudy (anonymous):

amen @BVB4rmy

OpenStudy (anonymous):

nmp

OpenStudy (anonymous):

So 58^2 = 24^2 + length^2 length^2= 3364 - 576 length^2 = 2788 length = 52.8015 52.8m

OpenStudy (anonymous):

I think o.o

OpenStudy (anonymous):

do you think thats the asnwer

OpenStudy (anonymous):

so now ur following me

OpenStudy (anonymous):

thx lol

OpenStudy (anonymous):

Yes I am 99% positive thats the answer ^~^

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

Your very welcome! :D